OpenStudy (anonymous):

The graph of which quadratic equation has a vertex of (3, 5)?

OpenStudy (amistre64):

the one that bends at 3,5 :)

OpenStudy (sriram):

(x-3)^2=y-5 or (y-5)^2=x-3

OpenStudy (sriram):

concept of shifting of origin